### v2

# Quasielectrostatic eigenmode analysis of anisotropic particles Asaf Farhi1 and Haim Suchowski1,21

Abstract
Anisotropic subwavelength particles combine the strong response and tunability of nanostructures with the exotic properties of anisotropic materials. Here we analytically derive the quasielectrostatic eigenmodes and eigenpermittivity relations of anistropic spheres and ellipsoids and present the results for spheres. We show that anisotropic particles exhibit novel radiation patterns and axial-eigenpermittivity sum rules, resulting in resonance splitting and eigenvalue degeneracy. Our results pave the way for a new generation of directional and multispectral detectors and biomarkers and apply to other fields of physics.
Subwavelength structures have attracted significant attention in various fields of physics including photonics, magnetisim, and quantum information, mainly due to their ability
to provide precise control over the spatial response
and their unique physical properties. A key distinction of subwavelength particles is their strong response, which can occur without gain [1, 2, 3, 4]. Furthermore, by adjusting their geometry, it is often possible to tailor their response at will. In photonics, they have been employed to manipulate waves in unprecedented ways, enabling to realize phenomena such as negative refractive index and spasing, as well as advancing applications in sensing, nonlinear emission, and cancer phototherapy [5, 6, 7, 8, 9, 10, 11]. While isotropic nanostructures have widespread use [12], anisotropic nanostructures are currently at the forefront of photonics research [13, 14].
When particles are composed of anisotropic media, they typically display resonances in unexplored spectral regions and exhibit exotic properties that arise from the directional and multispectral nature of the bulk [15], holding promise for a new generation of anisotropic resonators and biomarkers.
In addition to this class of resonators, a variety of naturally occurring and existing systems are, in fact, anisotropic particles, ranging from liquid crystal droplets to ferromagnetic particles and ice grains [16, 17, 18, 19, 20, 21]. Moreover, due to the correspondence between the optical properties of isotropic nanoparticles and atoms, better understanding of anisotropic particles may have implications for anisotropic molecules. [22, 23, 24, 25].
When the particle size is much smaller than the wavelength, a quasistatic analysis can be employed. In this regime,
Laplace’s equation plays a dominant role in a variety of fields including quasi-electrostatics, gravity, thermal physics, fluid mechanics, and magnetostatics [26, 27, 28, 29, 30, 31, 32]. Analytic solutions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [33, 34, 2, 35, 36]. Recently, the emergence of anisotropic materials in optics [13, 37, 38, 14, 39] has motivated the investigation of the response of anisotropic inclusions, with works on isotropic-mode-based scattering analysis from a uniaxial sphere, semianalytical and numerical analyses of a uniaxial spheroid, and eigenmode analysis of a biaxial slab [40, 4, 41, 42, 43]. More recently, an experiment on the modes and resonances of biaxial ellipsoids was performed for the first time by the authors and their colleagues [44].
However, while the dipole response of anisotropic particles is well understood [45], the derivation of higher-order modes and eigenpermittivities remains unexplored due to the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems.
The high-order response of such particles is important in two common scenarios: when the source is in proximity to the particle [14, 46] and when a group of close particles interact e.g., due to a far field excitation, and significantly enhance the field [47, 48]. Hence, addressing the higher-order response is crucial for unlocking this new generation of high-Q tunable, directional, and multispectral resonators and biomarkers that cover unexplored spectral regions, with a variety of novel applications in biomedicine, metamaterials, and photonics [14, 49, 50, 51, 44].
Figure 1: Angular distributions of the quasi-electrostatic potential eigenstates for isotropic, uniaxial, and biaxial spheres. The modes of the isotropic sphere are directed along the field/dipole excitation direction [1, 2] whereas the modes of the anisotropic spheres are aligned with the crystal axes and multispectral (colors correspond to frequencies). The second-order modes of anisotropic particles exhibit resonance splitting, and degeneracy for uniaxial spheres, leading to three superimposed modes.
Here we derive the analytic electroquasistatic eigenmodes and eigenpermittivity relations of anisotropic spheres and ellipsoids embedded in an isotropic medium.
While eigenvalue problems usually enable only one degree of freedom of the eigenvalue, here we utilize the definition of an eigenfunction that exists without a source to allow several degrees of freedom of the permittivities in the different axes. Our analysis applies to uniaxial and biaxial particles of any material and size as long as the quasistatic approximation holds. In addition, it can readily include the temporal dependence of the response by incorporating the time varying function of the excitation source. In our analysis we show that anisotropic particles exhibit novel radiation patterns. In addition, we identify axial-eigenpermittivity sum
rules, resulting in resonance splitting in anisotropic particles. We also show that the physics of anisotropic spheres is different from anisotropic ellipsoids: anisotropic spheres support uniaxial modes and exhibit eigenfrequency degeneracy, which leads to coherently superimposed modes, that anisotropic ellipsoids do not support. We apply our theory to spherical uniaxial and biaxial particles composed of hexagonal boron nitride (hBN) and molybdenum trioxide MoO3) phonon supporting polar crystals, which can be synthesized, and present their unique spectra and eigenmodes [3, 52, 53, 54].
Our starting point in calculating the eigenstates in the quasistatic regime is
solving Laplace’s equation without a source, that reads for
anisotropic media [55]
where is an electric potential eigenstate that exists without a source for the eigenpermittivity tensor
inside the inclusion and
in the host medium ( is the identity matrix). Note that the physical permittivity tensor generally depends on and the resonances occur when it approximately satisfies the eigenpermittivity relations. Clearly, have to satisfy
the boundary conditions at the particle interface. In addition to the challenge that the symmetry of
the material/crystal differs from the symmetry of the inclusion, Laplace’s equation has to be satisfied for anisotropic media, namely
As a result, the widely-employed spherical harmonics cannot be eigenstates
of anisotropic spherical particles.
In Fig. 1, left column, we present isotropic, uniaxial, and biaxial particles, which are the most common types of particles.
We now analyze the eigenmodes and eigenpermittivities of anisotropic
spheres and ellipsoids. We start with the dipole response
and proceed to the high-order mode responses.
The anisotropic
dipole eigenstates and eigenpermittivities of a sphere and ellipsoid
can be deduced from the scattering analyses in the literature [45, 36]
and read for a dipole mode oriented along
where in the case of an ellipsoid, are its intersections
with the positive axes, respectively, and the second expression for the ellipsoid is on the external envelope. While these eigenstates
and eigenpermittivities appear to be identical to the isotropic ones,
here the modes are oriented along the resonant crystal axis independently
of the incoming field polarization, even though the polarization direction
does affect the excitation magnitude.
We now proceed to the high-order modes of anisotropic particles. In our analysis we choose to express the
eigenstates in cartesian coordinates due to the crystal symmetry with functions
that are suitable for anisotropic medium i.e., ones that enable when substituted in Laplace’s equation and satisfy boundary conditions.
Since the field outside the sphere can be expanded with spherical harmonics,
the high-order spherical harmonics are suitable only for isotropic
medium [2, 35], and there
is continuity of the potential on the particle envelope, we suggest
that the anisotropic modes are composed of specific superpositions of isotropic modes.
Based on these considerations, we propose the following first second-order anisotropic sphere modes
where are the isotropic-sphere eigenstates [1, 2]. Inside the sphere
we write
where in order to normalize the modes one has to multiply them by
Imposing Laplace’s equation we obtain
which corresponds to a uniaxial medium inside the sphere.
is continuous at the interface by construction
and we impose continuity of We first calculate inside
the sphere by projecting
where we have used Note that the dot
product with retrieved that were removed in the differentiation,
which gives a factor of 2 as in the isotropic case, only that there
is a freedom of choice of
We then calculate outside the sphere
By equating them we obtain
Note that the resonance condition depends on which can be utilized to sense the environment in which the particle is situated.
Figure 2: Spectrum and eigenmodes of a subwavelength uniaxial sphere. (a) A uniaxial sphere and the permittivity of bulk hBN. (b) The scattering spectrum and of the dipole modes on an hBN sphere surface. Here we see scattering peaks at approximately (red line) and (blue line) associated with dipole modes in and respectively. The crystal axes [100],[010],[001] are directed along the axes, respectively. (c) The scattering spectrum and of the second-order modes, where the and modes and the uniaxial and modes are both doubly degenerate in frequency and there is a resonance splitting compared to the bulk material. (d) Scattering spectrum and of the third-order mode, where is excited at two s. In addition, there are geometric-anisotropic shifts compared to the bulk for all modes. Interestingly, high-order resonances have lower frequencies compared to the dipole modes, unlike isotropic spheres.
We now suggest the second second-order mode for an anisotropic sphere
and write the potential and field inside the sphere
where to normalize the modes one has to multiply them by Laplace’s equation is satisfied for this mode
We calculate at the boundaries
By equating these expressions we get
which for reads
Since the choice of system of coordinates is arbitrary, the derived modes
and eigenpermittivity relations should hold for any two axes. The
expression above is an interesting and novel eigenpermittivity sum rule,
this time for a general biaxial medium. Thus, for uniaxial and biaxial spheres four and three second-order modes can be excited, respectively. In addition, one can see that there is a degeneracy
for since both anisotropic
modes are excited. Coherently adding these modes gives for the angular part, where we assumed equal weights of the modes. Moreover, there is an additional degeneracy for a uniaxial sphere arising from the sum rule, and coherently adding the corresponding modes for gives Finally, for a resonance at a given axis, the sum rule is typically approached twice with the two other axes, leading to resonance splitting.
In Fig. 1 we present a scheme of the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles. It can be seen that anisotropic particles support significantly different modes, which appear in several spectral regions and their direction is determined only by the crystal orientation. Moreover, anisotropic particles exhibit rich physics with resonance splitting for both uniaxial and biaxial particles and degeneracy in frequency for uniaxial spheres, which leads to three superimposed modes.
Figure 3: Spectrum and eigenmodes of a subwavelength biaxial particle. (a) A biaxial sphere
with the bulk permittivities of . The scattering spectra and of the eigenmodes on a sphere surface for the dipole (b), second-order (c), and third-order (d) modes. Note that there are geometric-anisotropic frequency shifts for all modes and resonance splitting for the second-order modes compared to the bulk.
We now proceed to the third-order mode of an anisotropic sphere. We continue along the lines of the second-order biaxial mode and write Using the procedure above it can be shown that is the resonance condition. In Fig. 2 (a) we present a uniaxial sphere and the permittivities of a bulk composed of hBN with two resonances. We then show for a uniaxial sphere composed of hBN the spectrum and of the modes on the external surface of the sphere for the dipole (b), second-order (c), and third-order (d) modes. Interestingly, the particle exhibits geometric-anisotropic frequency shifts and the bulk resonance is split for the second-order modes, which also display a degeneracy.
In Fig. 3 we present a biaxial sphere and the axial permittivities of bulk with three resonances. We then show the scattering spectra and on the sphere surface for the dipole (b), second-order (c), and third-order modes (d). Similarly to the uniaxial particle, the second-order modes exhibit resonance splitting and all the modes display geometric-anisotropic frequency shifts. However, here the degeneracy is lifted since there are three different permittivities and the uniaxial modes are not present. Interestingly, in both uniaxial and biaxial spheres, the high-order mode resonances can have lower frequencies compared to the dipole mode, unlike the situation in isotropic spheres.
We now consider the second-order modes of an anisotropic ellipsoid in analogy to the anisotropic sphere.
We utilize inside the anisotropic
ellipsoid, similarly to the isotropic case of an ellipsoid in Ref. [36].
To impose continuity of we project in the
direction perpendicular to the interface with the unit vector
Comparing inside the ellipsoid to the isotropic case
we obtain that since and (inside and outside) are the same in both
cases, we have the general sum rule
and for the sphere limit we get as expected. The modes in the form similarly to the uniaxial sphere modes, have eigenpermittivity relations that are too restrictive and not likely to be realized experimentally.
We proceed to the third-order mode of an anisotropic ellipsoid. We write the potential and field inside the inclusion
Similarly, we compare between inside the ellipsoid in the isotropic and anisotropic cases and obtain that a resonance occurs when
We take the anisotropic-sphere limit of the eigenpermittivity relation and obtain agreement
We verified the theory for anisotropic ellipsoids with near-field spectroscopy of two nanoellipsoids composed of in Ref. [44] and obtained excellent agreement for both the modes and the spectra. In that work the near-field imaging tip scattered the field component in the axial direction to the far field, which was then measured.
In conclusion, we presented a framework to derive the
modes and eigenpermittivity relations of anisotropic spherical and
ellipsoid particles in the quasistatic regime. While isotropic structures
have discrete eigenpermittivities, we showed that anisotropic particles exhibit eigenpermittivity sum rules that couple
the axial permittivities. We identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes, and resonance splitting of the second-order modes of anisotropic particles. Our analysis revealed that both particle types support biaxial modes but anisotropic nanospheres support uniaxial modes that anisotropic ellipsoids typically do not support.
The multispectral behavior of anisotropic particles may allow them to function as a biomarker with enhanced specificity of the spectral signature. In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection [49].
Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [50, 56, 57]. Furthermore, our derived modes of anisotropic particles are expected to have implications for the radiation patterns in quantum transitions of atoms in magnetic fields and anisotropic molecules. Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing. Our results
apply also to other fields of physics such as heat conduction and quasimagnetostatics [46].
Acknowledgement
E. E. Narimanov, J. Jeffet, and Y. Reches are acknowledged for the insightful comments.
We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.

### v3

@@
 Anisotropic subwavelength particles combine the strong response and tunability of nanostructures with the exotic properties of anisotropic materials.
-Here we analytically derive the quasielectrostatic eigenmodes and eigenpermittivity relations of anistropic spheres and ellipsoids and present the results for spheres.
+Here we analytically derive the quasielectrostatic eigenmodes and eigenpermittivity relations of anistropic spheres and ellipsoids, and present the results for spheres.
 We show that anisotropic particles exhibit novel radiation patterns and axial-eigenpermittivity sum rules, resulting in resonance splitting and eigenvalue degeneracy.
@@
 In photonics, they have been employed to manipulate waves in unprecedented ways, enabling to realize phenomena such as negative refractive index and spasing, as well as advancing applications in sensing, nonlinear emission, and cancer phototherapy [5, 6, 7, 8, 9, 10, 11].
-While isotropic nanostructures have widespread use [12], anisotropic nanostructures are currently at the forefront of photonics research [13, 14].
+While isotropic nanostructures have widespread use [12], anisotropic nanostructures, predominantly composed of uniaxial and more recently biaxial materials, are currently at the forefront of photonics research [13, 14].
 When particles are composed of anisotropic media, they typically display resonances in unexplored spectral regions and exhibit exotic properties that arise from the directional and multispectral nature of the bulk [15], holding promise for a new generation of anisotropic resonators and biomarkers.
-In addition to this class of resonators, a variety of naturally occurring and existing systems are, in fact, anisotropic particles, ranging from liquid crystal droplets to ferromagnetic particles and ice grains [16, 17, 18, 19, 20, 21].
+In addition to this class of resonators, a variety of naturally occurring and existing systems are in fact uniaxial particles, ranging from liquid crystal droplets to ferromagnets and ice grains [16, 17, 18, 19, 20, 21].
 Moreover, due to the correspondence between the optical properties of isotropic nanoparticles and atoms, better understanding of anisotropic particles may have implications for anisotropic molecules.
@@
 Laplace’s equation plays a dominant role in a variety of fields including quasi-electrostatics, gravity, thermal physics, fluid mechanics, and magnetostatics [26, 27, 28, 29, 30, 31, 32].
-Analytic solutions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [33, 34, 2, 35, 36].
-Recently, the emergence of anisotropic materials in optics [13, 37, 38, 14, 39] has motivated the investigation of the response of anisotropic inclusions, with works on isotropic-mode-based scattering analysis from a uniaxial sphere, semianalytical and numerical analyses of a uniaxial spheroid, and eigenmode analysis of a biaxial slab [40, 4, 41, 42, 43].
-More recently, an experiment on the modes and resonances of biaxial ellipsoids was performed for the first time by the authors and their colleagues [44].
-However, while the dipole response of anisotropic particles is well understood [45], the derivation of higher-order modes and eigenpermittivities remains unexplored due to the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems.
+Analytic solutions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [33, 34, 2, 35, 36, 37, 38, 39, 40, 41, 42, 43].
+Recently, the emergence of anisotropic materials in optics [13, 44, 45, 14, 46] has motivated the investigation of the response of anisotropic inclusions, with semianalytical and numerical analyses of spheres and ellipsoids with cartesian anisotropy, eigenmode analysis of a biaxial slab, and studies of spheres with anisotropy in the radial direction [47, 4, 48, 49, 50, 51, 51, 52, 53, 54].
+More specifically, uniaxial and biaxial materials were shown to be very common in semiconductors and crystals, including: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium
+oxide (VO2), hexagonal boron nitride (hBN), and molybdenum trioxide MoO3) [55, 14, 56, 20].
+Very recently, an experiment on the modes and resonances of biaxial ellipsoids was performed for the first time by the authors and their colleagues [56].
+However, while the dipole response of anisotropic particles is well understood [42], the derivation of higher-order modes and eigenpermittivity relations remains unexplored due to the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems.
 The high-order response of such particles is important in two common scenarios: when the source is in proximity to the particle [14, 57] and when a group of close particles interact e.g., due to a far field excitation, and significantly enhance the field [58, 59].
@@
 The second-order modes of anisotropic particles exhibit resonance splitting, and degeneracy for uniaxial spheres, leading to three superimposed modes.
-Here we derive the analytic electroquasistatic eigenmodes and eigenpermittivity relations of anisotropic spheres and ellipsoids embedded in an isotropic medium.
-While eigenvalue problems usually enable only one degree of freedom of the eigenvalue, here we utilize the definition of an eigenfunction that exists without a source to allow several degrees of freedom of the permittivities in the different axes.
+Here we derive the analytic electroquasistatic eigenmodes and eigenpermittivities relations of anisotropic spheres and ellipsoids embedded in an isotropic medium.
+While eigenvalue problems usually enable only one degree of freedom of the eigenvalue, here we utilize the definition of an eigenfunction that exists without a source [1, 2], to allow several degrees of freedom of the permittivities in the different axes.
 Our analysis applies to uniaxial and biaxial particles of any material and size as long as the quasistatic approximation holds.
-In addition, it can readily include the temporal dependence of the response by incorporating the time varying function of the excitation source.
-In our analysis we show that anisotropic particles exhibit novel radiation patterns.
+In addition, it can readily include the temporal dependence of the response by incorporating the time-varying function of the excitation source.
+Note that whereas previous (semianalytical) works on particles with cartesian anisotropy focused on dielectric materials with nonresonant excitations, we consider the resonant modes with at least one [52, 51].
+Our analysis shows that anisotropic particles exhibit novel radiation patterns.
+While we consider the localized modes, in practice the particles have size and radiate and there is a one-to-one mapping between the near and far field modes, which determines the radiation patterns [63].
 In addition, we identify axial-eigenpermittivity sum
@@
 We also show that the physics of anisotropic spheres is different from anisotropic ellipsoids: anisotropic spheres support uniaxial modes and exhibit eigenfrequency degeneracy, which leads to coherently superimposed modes, that anisotropic ellipsoids do not support.
-We apply our theory to spherical uniaxial and biaxial particles composed of hexagonal boron nitride (hBN) and molybdenum trioxide MoO3) phonon supporting polar crystals, which can be synthesized, and present their unique spectra and eigenmodes [3, 52, 53, 54].
+We apply our theory to spherical uniaxial and biaxial particles composed of hBN and MoO3 phonon supporting polar crystals, which can be synthesized, and present their unique spectra and eigenmodes [3, 64, 65, 66].
 Our starting point in calculating the eigenstates in the quasistatic regime is
@@
 Clearly, have to satisfy
-the boundary conditions at the particle interface. In addition to the challenge that the symmetry of
+the boundary conditions at the particle interface.
+Once the source-free Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system.
+In addition to the challenge that the symmetry of
 the material/crystal differs from the symmetry of the inclusion, Laplace’s equation has to be satisfied for anisotropic media, namely
-As a result, the widely-employed spherical harmonics cannot be eigenstates
+As a result, the widely-employed spherical harmonics are in general not eigenstates
 of anisotropic spherical particles.
@@
 Since the field outside the sphere can be expanded with spherical harmonics,
-the high-order spherical harmonics are suitable only for isotropic
-medium [2, 35], and there
+the high-order spherical harmonics are valid for isotropic
+medium [2, 35],
+and there
 is continuity of the potential on the particle envelope, we suggest
-that the anisotropic modes are composed of specific superpositions of isotropic modes.
-Based on these considerations, we propose the following first second-order anisotropic sphere modes
-where are the isotropic-sphere eigenstates [1, 2]. Inside the sphere
-we write
-where in order to normalize the modes one has to multiply them by
-Imposing Laplace’s equation we obtain
-which corresponds to a uniaxial medium inside the sphere.
-is continuous at the interface by construction
-and we impose continuity of We first calculate inside
-the sphere by projecting
-where we have used Note that the dot
-product with retrieved that were removed in the differentiation,
-which gives a factor of 2 as in the isotropic case, only that there
-is a freedom of choice of
-We then calculate outside the sphere
-By equating them we obtain
-Note that the resonance condition depends on which can be utilized to sense the environment in which the particle is situated.
+that the anisotropic sphere modes are composed of specific isotropic-sphere modes or their superpositions [1, 2].
+To obtain eigenstates of a uniaxial sphere, we aim to arrive from anisotropic Laplace’s equation at without having to satisfy so that the modes can be excited when While the isotropic sphere eigenmodes and their superpositions satisfy anisotropic Laplace’s equation for when and it is sufficient to require .
+Interestingly, such functions can be composed of the isotropic-sphere eigenstates Thus, we find the uniaxial sphere modes and obtain their eigenpermittivity relations from the boundary conditions to arbitrary order:
+It can be seen that since in the boundary conditions and modes composed of and satisfy the isotropic eigenpermittivities [1, 2] for the equal permittivity axes and new eigenpermittivity relations, respectively.
+However, our modes in both cases are directional as opposed to the isotropic sphere modes.
+Note that the resonance conditions depend on which can be utilized to sense the environment in which the particle is situated.
 Figure 2: Spectrum and eigenmodes of a subwavelength uniaxial sphere.
@@
 Here we see scattering peaks at approximately (red line) and (blue line) associated with dipole modes in and respectively.
-The crystal axes [100],[010],[001] are directed along the axes, respectively.
+The crystal axes [100],[010],[001] are directed alfong the axes, respectively.
 (c) The scattering spectrum and of the second-order modes, where the and modes and the uniaxial and modes are both doubly degenerate in frequency and there is a resonance splitting compared to the bulk material.
@@
 Interestingly, high-order resonances have lower frequencies compared to the dipole modes, unlike isotropic spheres.
-We now suggest the second second-order mode for an anisotropic sphere
-and write the potential and field inside the sphere
-where to normalize the modes one has to multiply them by Laplace’s equation is satisfied for this mode
-We calculate at the boundaries
-By equating these expressions we get
-which for reads
-Since the choice of system of coordinates is arbitrary, the derived modes
-and eigenpermittivity relations should hold for any two axes. The
+In some cases superpositions of the isotropic sphere modes satisfy and anisotropic Laplace’s equation is valid for arbitrary values of
+Thus, we suggest the second-order biaxial-sphere modes and calculate the eigenpermittivity relations from the boundary condition:
+The
 expression above is an interesting and novel eigenpermittivity sum rule,
-this time for a general biaxial medium. Thus, for uniaxial and biaxial spheres four and three second-order modes can be excited, respectively. In addition, one can see that there is a degeneracy
-for since both anisotropic
-modes are excited.
-Coherently adding these modes gives for the angular part, where we assumed equal weights of the modes.
-Moreover, there is an additional degeneracy for a uniaxial sphere arising from the sum rule, and coherently adding the corresponding modes for gives Finally, for a resonance at a given axis, the sum rule is typically approached twice with the two other axes, leading to resonance splitting.
+this time for a general biaxial medium.
+In addition, one can see that there is a degeneracy
+for uniaxial particles with since both are excited.
+Note that one could select the basis for the uniaxial modes.
 In Fig.
 1 we present a scheme of the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles.
-It can be seen that anisotropic particles support significantly different modes, which appear in several spectral regions and their direction is determined only by the crystal orientation.
-Moreover, anisotropic particles exhibit rich physics with resonance splitting for both uniaxial and biaxial particles and degeneracy in frequency for uniaxial spheres, which leads to three superimposed modes.
+It can be seen that anisotropic particles support significantly different modes, which appear in several spectral regions, and their direction is determined only by the crystal orientation.
+Moreover, anisotropic particles exhibit rich physics with resonance splitting for both uniaxial and biaxial particles and degeneracy in frequency for uniaxial particles, which leads to superimposed modes.
 Figure 3: Spectrum and eigenmodes of a subwavelength biaxial particle. (a) A biaxial sphere
@@
 Note that there are geometric-anisotropic frequency shifts for all modes and resonance splitting for the second-order modes compared to the bulk.
-We now proceed to the third-order mode of an anisotropic sphere.
+We now proceed to the third-order biaxial mode of an anisotropic sphere.
 We continue along the lines of the second-order biaxial mode and write Using the procedure above it can be shown that is the resonance condition.
@@
 Interestingly, in both uniaxial and biaxial spheres, the high-order mode resonances can have lower frequencies compared to the dipole mode, unlike the situation in isotropic spheres.
-We now consider the second-order modes of an anisotropic ellipsoid in analogy to the anisotropic sphere.
+We now consider the second-order modes for anisotropic ellipsoids in analogy to the anisotropic sphere.
 We utilize inside the anisotropic
@@
 We take the anisotropic-sphere limit of the eigenpermittivity relation and obtain agreement
-We verified the theory for anisotropic ellipsoids with near-field spectroscopy of two nanoellipsoids composed of in Ref.
+We verified our theory for two biaxial ellipsoids composed of with hyperspectral midIR near-field imaging in Ref.
 [56] and obtained excellent agreement for both the modes and the spectra.
@@
 the axial permittivities.
-We identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes, and resonance splitting of the second-order modes of anisotropic particles.
+We identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes, and resonance splitting of anisotropic particles.
 Our analysis revealed that both particle types support biaxial modes but anisotropic nanospheres support uniaxial modes that anisotropic ellipsoids typically do not support.

### v4

@@
 Anisotropic subwavelength particles combine the strong response and tunability of nanostructures with the exotic properties of anisotropic materials.
-Here we analytically derive the quasielectrostatic eigenmodes and eigenpermittivity relations of anistropic spheres and ellipsoids, and present the results for spheres.
-We show that anisotropic particles exhibit novel radiation patterns and axial-eigenpermittivity sum rules, resulting in resonance splitting and eigenvalue degeneracy.
+Here we derive closed-form expressions for the quasielectrostatic eigenmodes and eigenpermittivity relations of uniaxial and biaxial spheres and ellipsoids.
+We show that such particles exhibit novel radiation patterns and axial-eigenpermittivity sum rules, resulting in resonance splitting and eigenvalue degeneracy.
+Finally, we analytically predict the Q-factors of the resonance peaks.
 Our results pave the way for a new generation of directional and multispectral detectors and biomarkers and apply to other fields of physics.
@@
 When the particle size is much smaller than the wavelength, a quasistatic analysis can be employed. In this regime,
-Laplace’s equation plays a dominant role in a variety of fields including quasi-electrostatics, gravity, thermal physics, fluid mechanics, and magnetostatics [26, 27, 28, 29, 30, 31, 32].
-Analytic solutions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [33, 34, 2, 35, 36, 37, 38, 39, 40, 41, 42, 43].
+Laplace’s equation plays a dominant role in a variety of fields including electrodynamics, gravity, thermal physics, fluid mechanics, and magnetism [26, 27, 28, 29, 30, 31, 32].
+Closed-form resonance conditions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [33, 34, 2, 35, 36, 37, 38, 39, 40, 41, 42, 43].
 Recently, the emergence of anisotropic materials in optics [13, 44, 45, 14, 46] has motivated the investigation of the response of anisotropic inclusions, with semianalytical and numerical analyses of spheres and ellipsoids with cartesian anisotropy, eigenmode analysis of a biaxial slab, and studies of spheres with anisotropy in the radial direction [47, 4, 48, 49, 50, 51, 51, 52, 53, 54].
 More specifically, uniaxial and biaxial materials were shown to be very common in semiconductors and crystals, including: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium
-oxide (VO2), hexagonal boron nitride (hBN), and molybdenum trioxide MoO3) [55, 14, 56, 20].
+oxide (VO2), hexagonal boron nitride (hBN), and molybdenum trioxide MoO3,) and various geometries have been investigated [55, 14, 56, 20].
 Very recently, an experiment on the modes and resonances of biaxial ellipsoids was performed for the first time by the authors and their colleagues [56].
-However, while the dipole response of anisotropic particles is well understood [42], the derivation of higher-order modes and eigenpermittivity relations remains unexplored due to the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems.
-The high-order response of such particles is important in two common scenarios: when the source is in proximity to the particle [14, 57] and when a group of close particles interact e.g., due to a far field excitation, and significantly enhance the field [58, 59].
+However, while previous approaches have successfully addressed particles composed of isotropic materials by deriving closed-form expressions of the eigenstates and resonance conditions, fundamental challenges are encountered when analyzing anisotropic systems, particularly in predicting higher-order mode behavior.
+These challenges arise from the mismatch between the particles’ geometric and crystal symmetries, along with the additional degrees of freedom introduced by the axial permittivities, which differentiates it from standard eigenvalue problems.
+This motivates our exploration of a new theoretical framework.
+The high-order response of anisotropic particles is important in two common scenarios: when the source is in proximity to the particle [14, 57] and when a group of close particles interact e.g., due to a far field excitation, and significantly enhance the field [58, 59].
 Hence, addressing the higher-order response is crucial for unlocking this new generation of high-Q tunable, directional, and multispectral resonators and biomarkers that cover unexplored spectral regions, with a variety of novel applications in biomedicine, metamaterials, and photonics [14, 60, 61, 62, 56].
@@
 The modes of the isotropic sphere are directed along the field/dipole excitation direction [1, 2] whereas the modes of the anisotropic spheres are aligned with the crystal axes and multispectral (colors correspond to frequencies).
-The second-order modes of anisotropic particles exhibit resonance splitting, and degeneracy for uniaxial spheres, leading to three superimposed modes.
+The high-order modes of anisotropic particles exhibit resonance splitting, and degeneracy for uniaxial spheres, leading to superimposed modes.
 Here we derive the analytic electroquasistatic eigenmodes and eigenpermittivities relations of anisotropic spheres and ellipsoids embedded in an isotropic medium.
@@
 In addition, it can readily include the temporal dependence of the response by incorporating the time-varying function of the excitation source.
-Note that whereas previous (semianalytical) works on particles with cartesian anisotropy focused on dielectric materials with nonresonant excitations, we consider the resonant modes with at least one [52, 51].
+Note that whereas previous (semianalytical) works on particles with cartesian anisotropy primarly focused on dielectric materials with all axial permittivities positive [63, 51], here we explore the resonant modes that emerge when at least one of the principal components is negative .
 Our analysis shows that anisotropic particles exhibit novel radiation patterns.
-While we consider the localized modes, in practice the particles have size and radiate and there is a one-to-one mapping between the near and far field modes, which determines the radiation patterns [63].
+While we consider the localized modes, in practice subwavelength particles radiate and there is a one-to-one mapping between the near and far field modes, which determines the radiation patterns, see Supplementary Material (SM) for details.
+Even though the radiation of subwavelength particles in free space via high-order modes is negligible, when there is one or more particles or excitation source at a distance from the particle, these modes play a more significant role in the radiation.
 In addition, we identify axial-eigenpermittivity sum
@@
 the boundary conditions at the particle interface.
-Once the source-free Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system.
+Once the source-free Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system by definition.
 In addition to the challenge that the symmetry of
@@
 dipole eigenstates and eigenpermittivities of a sphere and ellipsoid
-can be deduced from the scattering analyses in the literature [42, 36]
+can be deduced from the scattering analyses in the literature [42]
 and read for a dipole mode oriented along
@@
 with the positive axes, respectively, and the second expression for the ellipsoid is on the external envelope. While these eigenstates
-and eigenpermittivities appear to be identical to the isotropic ones,
+and eigenpermittivities appear to be identical to the isotropic ones [36],
 here the modes are oriented along the resonant crystal axis independently
-of the incoming field polarization, even though the polarization direction
+of the incoming field polarization, even though it
 does affect the excitation magnitude.
@@
 that the anisotropic sphere modes are composed of specific isotropic-sphere modes or their superpositions [1, 2].
-To obtain eigenstates of a uniaxial sphere, we aim to arrive from anisotropic Laplace’s equation at without having to satisfy so that the modes can be excited when While the isotropic sphere eigenmodes and their superpositions satisfy anisotropic Laplace’s equation for when and it is sufficient to require .
+To obtain eigenstates of a uniaxial sphere, we aim to arrive from anisotropic Laplace’s equation at without having to satisfy so that the modes can be excited when While the isotropic sphere eigenmodes and their superpositions satisfy anisotropic Laplace’s equation for when and it is sufficient to require to satisfy the equation.
 Interestingly, such functions can be composed of the isotropic-sphere eigenstates Thus, we find the uniaxial sphere modes and obtain their eigenpermittivity relations from the boundary conditions to arbitrary order:
@@
 However, our modes in both cases are directional as opposed to the isotropic sphere modes.
-Note that the resonance conditions depend on which can be utilized to sense the environment in which the particle is situated.
+Alternatively, one could select the basis .
+Note also that the resonance conditions depend on which can be utilized to sense the environment in which the particle is situated.
 Figure 2: Spectrum and eigenmodes of a subwavelength uniaxial sphere.
@@
 Here we see scattering peaks at approximately (red line) and (blue line) associated with dipole modes in and respectively.
-The crystal axes [100],[010],[001] are directed alfong the axes, respectively.
+The crystal axes [100],[010],[001] are directed along the axes, respectively.
 (c) The scattering spectrum and of the second-order modes, where the and modes and the uniaxial and modes are both doubly degenerate in frequency and there is a resonance splitting compared to the bulk material.
 (d) Scattering spectrum and of the third-order mode, where is excited at two s.
-In addition, there are geometric-anisotropic shifts compared to the bulk for all modes.
-Interestingly, high-order resonances have lower frequencies compared to the dipole modes, unlike isotropic spheres.
-In some cases superpositions of the isotropic sphere modes satisfy and anisotropic Laplace’s equation is valid for arbitrary values of
-Thus, we suggest the second-order biaxial-sphere modes and calculate the eigenpermittivity relations from the boundary condition:
-The
-expression above is an interesting and novel eigenpermittivity sum rule,
+Moreover, there are geometric-anisotropic shifts compared to the bulk for all modes.
+Interestingly, high-order resonances have lower compared to the dipole modes, unlike isotropic spheres.
+In some cases superpositions of the isotropic sphere modes satisfy and anisotropic Laplace’s equation is valid for arbitrary values of , which have to obey only the boundary condition.
+Thus, we suggest the following biaxial-sphere modes and calculate the eigenpermittivity relations from the boundary condition:
+Note that this class of functions preserves the form under Thus, the axial permittivities satisfy novel and interesting sum rules,
 this time for a general biaxial medium.
-In addition, one can see that there is a degeneracy
-for uniaxial particles with since both are excited.
-Note that one could select the basis for the uniaxial modes.
 In Fig.
-1 we present a scheme of the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles.
+1 we present the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles.
 It can be seen that anisotropic particles support significantly different modes, which appear in several spectral regions, and their direction is determined only by the crystal orientation.
@@
 Note that there are geometric-anisotropic frequency shifts for all modes and resonance splitting for the second-order modes compared to the bulk.
-We now proceed to the third-order biaxial mode of an anisotropic sphere.
-We continue along the lines of the second-order biaxial mode and write Using the procedure above it can be shown that is the resonance condition.
+Furthermore, for a biaxial material whose bulk exhibits three resonances, our resonance conditions predict 18 different peaks in the spectrum in addition to the 3 dipole resonances.
+In the SM we show that the derived uniaxial modes are all the modes and if additional biaxial modes exist they are typically negligible.
 In Fig.
@@
 We then show for a uniaxial sphere composed of hBN the spectrum and of the modes on the external surface of the sphere for the dipole (b), second-order (c), and third-order (d) modes.
-Interestingly, the particle exhibits geometric-anisotropic frequency shifts and the bulk resonance is split for the second-order modes, which also display a degeneracy.
+Interestingly, the particle exhibits geometric-anisotropic frequency shifts and the bulk resonance is split for the high-order modes, which also display a degeneracy.
 In Fig.
@@
 Interestingly, in both uniaxial and biaxial spheres, the high-order mode resonances can have lower frequencies compared to the dipole mode, unlike the situation in isotropic spheres.
-We now consider the second-order modes for anisotropic ellipsoids in analogy to the anisotropic sphere.
+We now consider the high-order modes of anisotropic ellipsoids in analogy to the anisotropic sphere.
+Interestingly, the biaxial-sphere modes inside the inclusion volume correspond to the second type of Niven’s functions used in isotropic ellipsoids, at least up to fourth order [36, 68, 69, 70]
 We utilize inside the anisotropic
 ellipsoid, similarly to the isotropic case of an ellipsoid in Ref. [36].
-To impose continuity of we project in the
+To impose continuity of for we project in the
 direction perpendicular to the interface with the unit vector
 Comparing inside the ellipsoid to the isotropic case
-we obtain that since and (inside and outside) are the same in both
-cases, we have the general sum rule
+we obtain that since and (inside and outside) are the same, we can equate to get an eigenmode and obtain the general sum rule
 and for the sphere limit we get as expected.
@@
 We proceed to the third-order mode of an anisotropic ellipsoid. We write the potential and field inside the inclusion
-Similarly, we compare between inside the ellipsoid in the isotropic and anisotropic cases and obtain that a resonance occurs when
+Similarly, we compare inside the ellipsoid in the isotropic and anisotropic cases and obtain that a resonance occurs when
 We take the anisotropic-sphere limit of the eigenpermittivity relation and obtain agreement
-We verified our theory for two biaxial ellipsoids composed of with hyperspectral midIR near-field imaging in Ref.
-[56] and obtained excellent agreement for both the modes and the spectra.
-In that work the near-field imaging tip scattered the field component in the axial direction to the far field, which was then measured.
+In this way one can also find the fourth-order biaxial ellipsoid mode and resonance condition.
+Finally, we verified our theory for uniaxial and biaxial spheres and a biaxial ellipsoid in simulations (see Appendix) and also experimentally for two biaxial ellipsoids with hyperspectral midIR near-field imaging in Ref.
+[56], both with excellent agreement for the modes and the spectra.
 In conclusion, we presented a framework to derive the
@@
 the axial permittivities.
-We identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes, and resonance splitting of anisotropic particles.
+We identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes, and resonance splitting in anisotropic particles.
 Our analysis revealed that both particle types support biaxial modes but anisotropic nanospheres support uniaxial modes that anisotropic ellipsoids typically do not support.
+Importantly, our derived modes can be used to calculate the eigenstates and resonance conditions of clusters of anisotropic particles [35].
 The multispectral behavior of anisotropic particles may allow them to function as a biomarker with enhanced specificity of the spectral signature.
@@
 Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing.
-Our results
+Finally, our results
 apply also to other fields of physics such as heat conduction and quasimagnetostatics [57].
 Acknowledgement
-E. E. Narimanov, J. Jeffet, and Y. Reches are acknowledged for the insightful comments.
+E. E. Narimanov, A. Niv, Y. Mazor, J. Jeffet, and Y. Reches are acknowledged for the insightful comments.
 We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.
+I Appendix
+Comparison of the theory to simulations
+Figure 4: Scattering spectrum, peak widths, and fields near resonance for a uniaxial hBN sphere of a radius excited by an oscillating dipole located at with a dipole moment of .
+Top: over the particle envelope calculated in COMSOL and compared to the peaks predicted analytically using with excellent agreement.
+The circles are the peak widths obtained analytically from compared to the ones calculated from the simulations, with very good agreement.
+(a),(b),(c),(d),(e) for the resonance conditions: calculated at
+respectively, using COMSOL compared to the mode fields calculated analytically with very good agreement.
+Bottom: over the particle envelope calculated in COMSOL for a dipole located at (same dipole distance as the previous case) with a dipole moment of As predicted the modes disappeared from the spectrum.Figure 5: Scattering spectrum and modes/fields near resonance for a biaxial -MoO3 sphere of a radius excited by an oscillating dipole located at with a dipole moment of .
+Top: calculated in COMSOL and compared to the peaks predicted analytically using with excellent agreement.
+(a),(b),(c),(d),(e),(f),(g) for the resonance conditions: calculated at respectively, using COMSOL compared to the mode fields calculated analytically with very good agreement.
+Figure 6: Scattering spectrum and modes/fields near resonance for a biaxial -MoO3 ellipsoid with excited by an oscillating dipole located at with a dipole moment of .
+Top: Surface integral over the particle envelope calculated in COMSOL and compared to the peaks predicted analytically using with excellent agreement.
+(a),(b),(c),(d),(e),(f),(g) for the resonance conditions: calculated at respectively, using COMSOL with very good agreement to the mode fields calculated analytically (same as in Fig.
+5, defined in the ellipsoid volume).
+We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes and the permittivity model used in Figs.
+2,3 [55].
+The sphere is of radius and we situate an oscillating dipole at with a dipole moment of We perform frequency domain simulations with a frequency sweep from (1/s) to (1/s) with 304 frequencies using extensive COMSOL simulations.
+The dipole moment was chosen to be in the direction of since we assume [67, 73, 2], and later on prove, that the mode excitation coefficient depends on so that we need to incorporate only preserving the angular dependence in .
+This, together with the fact that the dipole has only components, ensures that the degenerate modes will add constructively, see SM for details.
+Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflects the presence of higher spatial frequencies near the dipole, which excite higher-order modes.
+We then calculate the peak widths analytically and from the simulations.
+To demonstrate the mode coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the axis.
+In this case we do not expect the modes to show in the spectrum since in order to excite them a component of the dipole location is required.
+In Fig.
+4 we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of compared to the surface integral of over the uniaxial particle envelope calculated in COMSOL with excellent agreement.
+Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows in Fig.
+4) and mode degeneracy for uniaxial spheres.
+We also obtained the peak widths analytically from and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to is altered due to the peak on the right, and therefore we haven’t accounted for it).
+Clearly, there is remarkable agreement between them.
+This showcases our important ability to analytically predict the Q factors given by
+We then plot for all the observed peaks inside the particle calculated in COMSOL and obtained analytically from the modes.
+While the COMSOL fields include the incoming field from the oscillating dipole, close to a resonance this is typically negligible and indeed we observe very good agreement.
+Moreover, we plot for a dipole located at the same distance as the previous case and oriented along (31nm).
+As we predicted the modes disappeared from the spectrum, verifying the dependency of the mode excitation coefficient on the dipole location and orientation.
+This highlights the directional nature of the modes as opposed to isotropic particles.
+We then considered a biaxial subwavelength sphere composed of MoO3 with the permittivity model in Ref. [55]. The sphere is excited by a dipole located at with the dipole moment of
+We performed frequency domain simulations with a frequency sweep between and with 175 frequencies using COMSOL.
+Here, again, the dipole moment is directed along to preserve the angular dependency of the mode in the mode-excitation coefficient.
+In Fig.
+5 we present the result of the surface integral over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement.
+The predicted peaks were readily calculated by finding the maxima of .
+Crucially, our simulations confirm the resonance splitting effect predicted analytically, see arrows in Fig.
+5.
+We then compared for all the observed resonances from the simulations to the ones obtained analytically from the modes with excellent agreement.
+We also report very good agreement between the peak frequency widths obtained analytically and the simulations (accurate comparison would require a highly-dense frequency sweep and simulations that are very extensive).
+Finally, we simulated a biaxial subwavelength ellipsoid of the most general geometry with composed of MoO3 with the permittivity model in Ref.
+[55].
+The ellipsoid is excited by a dipole located at with the dipole moment of
+We performed frequency domain simulations with a frequency sweep between and with 175 frequencies using COMSOL.
+In Fig.
+6 we present the result of the surface integral over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement.
+The predicted peaks were easily calculated by finding the maxima of .
+We then compared for all the observed resonances from the simulations to the ones obtained analytically from the modes with excellent agreement.
+We note that the isotropic ellipsoid eigenpermittivities that were calculated to obtain the biaxial ellipsoid resonance permittivity relations are the following:
+As can be seen, for an isotropic ellipsoid with these semi axis ratios and and due to the anisotropic material these resonance are split, as can be seen in Fig.
+6.
+This differs from the case of spheres where the splitting occurs for modes of the same order.
+Also, in this case the anisotropy of the material could in principle lead to resonance merging.
+Finally, there is a clear physical intuition for the resonance conditions of anisotropic ellipsoids: the resonances of isotropic ellispoids, on which they are based, tend to favor shorter ellipsoid axes, since they exhibit higher real permittivity values that have smaller imaginary parts of the physical permittivity, leading to stronger and sharper resonances.
+II Supplementary Material
+II.1 Correspondence between the derived modes and the radiation patterns
+Here we derive the correspondence between the localized modes that we obtained in the main text and the radiation patterns of the anisotropic subwavelength spheres.
+We utilize the TM vector spherical harmonics (VSH).
+This field pattern is proportional to the sphere mode outside the inclusion [74], which gives the quasistatic results in the limit of Even though this mode is usually considered for isotropic spheres, the results can be applied to the modes of anisotropic spheres since their modes are superpositions of spherical harmonics, which correspond to superpositions of VSHs.
+We write a TM VSH:
+where is the spherical Hankel function of the first kind.
+Using the vector identity for
+we get for the first term:
+where
+For the second term we use the identity
+to get
+Then, we utilize the spherical-harmonic identity
+to obtain
+All in all we get:
+We see that when the second term, which is longitudinal, dominates with scaling, and for
+the first term, which is transverse, dominates (actually its first term that scales as
+Thus, one can conclude from the superposition of that participate in the
+quasistatic anisotropic modes, that the radiation pattern will be of the form:
+Please note that which gives
+overall scaling of the radiated field and we get the following radiation pattern:
+II.2 Biaxial sphere modes as superpositions of spherical harmonics
+We express the second-order biaxial modes as superpositions of spherical harmonics:
+We now proceed to the third and fourth-order biaxial modes.
+We continue along the lines of the second-order biaxial mode and write Using the procedure above, it can be shown that and are the resonance conditions, respectively.
+Note that the modes and their resonance conditions readily follow.
+II.3 Completeness of the solution
+We now prove that the derived uniaxial-sphere modes are all the eigenmodes.
+Let us assume that there is an additional eigenstate besides the four modes for each .
+Since the isotropic-sphere modes span the space inside the sphere, this eigenstate is a superposition of the remaining isotropic-sphere modes
+An eigenstate must satisfy anisotropic Laplace’s equation inside the inclusion, which reads:
+Since it is a superposition of isotropic-sphere modes we write:
+However, all these modes have orders of that are 2 or higher (including from ) and each function is fundamentally different from the others.
+Hence, such an eigenstate does not exist, and we have a contradiction.
+Conclusion: the derived uniaxial sphere modes are all the modes.
+Note that the uniaxial mode basis does not span space as in the standard case.
+Similarly, we write for a biaxial particle
+If we get which is a uniaxial medium, and we therefore exclude this case.
+All the modes are isotropic (from the boundary conditions) and the modes satisfy when Even if such modes exist, their axial permittivities must satisfy both Laplace’s equation (in a non-trivial manner) and the boundary condition, which imposes strong constraints.
+As a result, their potential contribution is typically weak.
+Therefore, we expect that the dominant biaxial modes satisfy and considered all such possible modes in the main text.
+II.4 Superimposing the degenerate uniaxial modes for the sphere simulation comparison
+II.4.1 Theory
+We now analytically derive the field distributions and resonance frequencies for the uniaxial sphere and the dipole location in the Appendix.
+Dipole mode
+The resonance condition is and we therefore find the maximum of
+which occurs for hBN [55] at The mode field inside the particle is given by
+2nd order modes
+For the first 2nd-order mode we get from the theory
+which for hBN approximately occurs at We add the degenerate modes, which are excited with equal strengths:
+second mode
+The derived resonance condition reads
+For hBN this is approximately satisfied for
+Now we calculate the field by adding the degenerate modes:
+3rd order modes
+First mode
+The resonance condition reads:
+For hBN this occurs at
+. Adding the degenerate modes we get
+To simplify things, we look at the plane
+Second mode
+The resonance condition is satisfied
+for hBN for Adding the modes we obtain:
+We plot these field magnitude distributions in Fig. 4.
+II.5 Eigenmode expansion for an anisotropic inclusion
+Eigenstate expansions have been derived for an isotropic inclusion and for an inclusion with only one axial permittivity that differs from the host medium permittivity [74, 67].
+Here we expand the quasielectrostatic potential for an anisotropic inclusion.
+For concreteness, we consider a uniaxial inclusion and the derived eigenstates. Generalizing it to a biaxial inclusion readily follows.
+We start by writing Poisson’s equation for a medium with an anisotropic inclusion:
+where is a step function that equals 1 inside the inclusion, is the inclusion permittivity tensor, and is the host medium permittivity.
+Now we move one term to the lhs in order to be able to define an
+eigenvalue equation.
+where . The Green function in this case will be for a medium with an inclusion
+with anisotropy in the axis and axial permittivities that are equal to the host medium permittivity in the other axes.
+where we defined to adjust it to an eigenvalue equation with an operator operating on and we performed integration by parts.
+We express the electric potential and expand it using the anisotropic
+eigenfunctions defined by Laplace’s equation. In the following is defined for the particle with anisotropy in the axis. We write:
+Now we expand using the eigenstates of a particle with anisotropy in the axis:
+where is the electric potential for the dipole in free space.
+All in all we get:
+One can express the eigenvalue in the denominator in terms of the physical parameter using the eigenpermittivity relations, obtaining exactly the resonance conditions now for the physical parameters used along the paper (e.g., ones used in Figs.
+2 and 3).
+Note that the first sum includes only the dipole term and the second sum includes the remaining eigenstates in agreement with our analysis along the paper.
+It is worth noting that this formalism should also apply to other types of anisotropy such as spherical.
+Note that in Fig.
+4 there is a very small peak at This peak may be due to an off-resonance uniaxial mode as could be understood from the above expansion when applied to biaxial inclusions.
+Such an expansion is composed of three sums where the second one corresponds to the off resonance uniaxial-mode contributions.
+This qualitatively agrees with the potential calculated from the field in the simulation results of the form similarly to the mode constructed by adding the modes.
+Interestingly, this very small peak is not present in the spectrum of the biaxial ellipsoid, in agreement with our prediction that ellipsoids do not favor uniaxial modes.

### v5

@@
-Quasielectrostatic eigenmode analysis of anisotropic particles Asaf Farhi1 and Haim Suchowski1,21
+Universal framework for anisotropic particles with resonance laws and splitting
 Abstract
-Anisotropic subwavelength particles combine the strong response and tunability of nanostructures with the exotic properties of anisotropic materials.
-Here we derive closed-form expressions for the quasielectrostatic eigenmodes and eigenpermittivity relations of uniaxial and biaxial spheres and ellipsoids.
-We show that such particles exhibit novel radiation patterns and axial-eigenpermittivity sum rules, resulting in resonance splitting and eigenvalue degeneracy.
-Finally, we analytically predict the Q-factors of the resonance peaks.
-Our results pave the way for a new generation of directional and multispectral detectors and biomarkers and apply to other fields of physics.
-Subwavelength structures have attracted significant attention in various fields of physics including photonics, magnetisim, and quantum information, mainly due to their ability
-to provide precise control over the spatial response
-and their unique physical properties.
-A key distinction of subwavelength particles is their strong response, which can occur without gain [1, 2, 3, 4].
-Furthermore, by adjusting their geometry, it is often possible to tailor their response at will.
-In photonics, they have been employed to manipulate waves in unprecedented ways, enabling to realize phenomena such as negative refractive index and spasing, as well as advancing applications in sensing, nonlinear emission, and cancer phototherapy [5, 6, 7, 8, 9, 10, 11].
+Nanophotonics enables unprecedented control over light-matter interactions, yet conventional isotropic materials limit the spectral range and richness of mode responses in subwavelength structures.
+Anisotropic subwavelength particles—ubiquitous in natural and engineered systems—offer new degrees of freedom that couple geometry and material properties, unlocking previously inaccessible spectral regions.
+Here, we establish a universal quasistatic framework describing the eigenmodes and resonance conditions of uniaxial and biaxial nanoparticles.
+Closed-form solutions reveal axial-permittivity sum rules and a fundamental material-anisotropy-induced symmetry breaking, manifesting as resonance splitting and novel radiation patterns.
+Generalizing the theory to ellipsoids provides geometric tunability of the multispectral response, while analytic predictions of quality factors elucidate how anisotropy governs mode localization and energy loss.
+Full-wave simulations of h-BN and -MoO3 nanoparticles confirm the theory and highlight its predictive power.
+This framework unifies the understanding of anisotropic nanostructures across optics, thermal transport, and magnetism, opening pathways to a new generation of photonic devices with tunable multispectral response and controlled emission patterns, including directional detectors and spectrally selective mid infrared sensors.
+Main
+Subwavelength structures have attracted significant attention in various fields of physics including photonics, magnetisim, and quantum information, mainly due to their ability to provide precise control over the spatial response In photonics, they have been employed to manipulate waves in unprecedented ways, enabling to realize phenomena such as negative refractive index and spasing, as well as advancing applications in sensing, nonlinear emission, and cancer phototherapy [1, 2, 3, 4, 5, 6, 7].
 While isotropic nanostructures have widespread use [8], anisotropic nanostructures, predominantly composed of uniaxial and more recently biaxial materials, are currently at the forefront of photonics research [9, 10].
@@
 In addition to this class of resonators, a variety of naturally occurring and existing systems are in fact uniaxial particles, ranging from liquid crystal droplets to ferromagnets and ice grains [12, 13, 14, 15, 16, 17].
-Moreover, due to the correspondence between the optical properties of isotropic nanoparticles and atoms, better understanding of anisotropic particles may have implications for anisotropic molecules.
-[22, 23, 24, 25].
-When the particle size is much smaller than the wavelength, a quasistatic analysis can be employed. In this regime,
-Laplace’s equation plays a dominant role in a variety of fields including electrodynamics, gravity, thermal physics, fluid mechanics, and magnetism [26, 27, 28, 29, 30, 31, 32].
-Closed-form resonance conditions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [33, 34, 2, 35, 36, 37, 38, 39, 40, 41, 42, 43].
+Moreover, due to the correspondence between the optical properties of isotropic nanoparticles and atoms, better understanding of anisotropic particles may have implications for anisotropic molecules [18, 19, 20, 21].
+When the particle size is much smaller than the wavelength, a quasistatic analysis can be employed.
+In this regime, Laplace’s equation plays a dominant role in a variety of fields including electrodynamics, gravity, thermal physics, fluid mechanics, and magnetism [22, 23, 24, 25, 26, 27, 28].
+Analytic resonance conditions and eigenstates of isotropic Laplace’s equation have been mainly derived for structures with a high degree of symmetry [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39].
+Remarkably, to date, the only structure in physics for which closed-form quasistatic eigenmodes and eigenvalues have been derived is the subwavelength isotropic sphere [40, 41, 31].
 Recently, the emergence of anisotropic materials in optics [9, 42, 43, 10, 44] has motivated the investigation of the response of anisotropic inclusions, with semianalytical and numerical analyses of spheres and ellipsoids with cartesian anisotropy, eigenmode analysis of a biaxial slab, and studies of spheres with anisotropy in the radial direction [45, 46, 47, 48, 49, 50, 50, 51, 52, 53].
-More specifically, uniaxial and biaxial materials were shown to be very common in semiconductors and crystals, including: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium
-oxide (VO2), hexagonal boron nitride (hBN), and molybdenum trioxide MoO3,) and various geometries have been investigated [55, 14, 56, 20].
-Very recently, an experiment on the modes and resonances of biaxial ellipsoids was performed for the first time by the authors and their colleagues [56].
+Moreover, uniaxial and biaxial materials have been shown to be very common in semiconductors and crystals, including the following: SiC (4H), AlN, GaN, quartz, sapphire, calcite, vanadium oxide (VO2), hexagonal boron nitride (hBN), and molybdenum trioxide MoO3), with various geometries investigated [54, 10, 55, 16].
+Very recently, an experiment on the modes and resonances of biaxial nanoparticles was performed for the first time by the authors and their colleagues [55].
 However, while previous approaches have successfully addressed particles composed of isotropic materials by deriving closed-form expressions of the eigenstates and resonance conditions, fundamental challenges are encountered when analyzing anisotropic systems, particularly in predicting higher-order mode behavior.
@@
 Hence, addressing the higher-order response is crucial for unlocking this new generation of high-Q tunable, directional, and multispectral resonators and biomarkers that cover unexplored spectral regions, with a variety of novel applications in biomedicine, metamaterials, and photonics [10, 59, 60, 61, 55].
-Figure 1: Angular distributions of the quasi-electrostatic potential eigenstates for isotropic, uniaxial, and biaxial spheres.
-The modes of the isotropic sphere are directed along the field/dipole excitation direction [1, 2] whereas the modes of the anisotropic spheres are aligned with the crystal axes and multispectral (colors correspond to frequencies).
-The high-order modes of anisotropic particles exhibit resonance splitting, and degeneracy for uniaxial spheres, leading to superimposed modes.
-Here we derive the analytic electroquasistatic eigenmodes and eigenpermittivities relations of anisotropic spheres and ellipsoids embedded in an isotropic medium.
+Here, we present the second-ever closed-form eigenmode and eigenpermittivity solution for a structure, derived for uniaxial and biaxial spheres embedded in an isotropic medium.
 While eigenvalue problems usually enable only one degree of freedom of the eigenvalue, here we utilize the definition of an eigenfunction that exists without a source [41, 31], to allow several degrees of freedom of the permittivities in the different axes.
+We identify axial-eigenpermittivity sum rules, leading to resonance splitting in anisotropic particles.
+In addition, our analysis shows that anisotropic spheres exhibit novel radiation patterns.
+Furthermore, we generalize our results to uniaxial and biaxial ellipsoids, obtaining the corresponding eigenpermittivity sum rules.
+Moreover, we analytically derive the Q factors for such anisotropic particles.
+We apply our theory to uniaxial and biaxial particles composed of hBN and MoO3 phonon supporting polar crystals, which can be synthesized, and present their unique spectra and eigenmodes [62, 63, 64, 65, 55].
+Finally, we compare our results to full-wave simulations and find excellent agreement.
 Our analysis applies to uniaxial and biaxial particles of any material and size as long as the quasistatic approximation holds.
@@
 Note that whereas previous (semianalytical) works on particles with cartesian anisotropy primarly focused on dielectric materials with all axial permittivities positive [66, 50], here we explore the resonant modes that emerge when at least one of the principal components is negative .
-Our analysis shows that anisotropic particles exhibit novel radiation patterns.
-While we consider the localized modes, in practice subwavelength particles radiate and there is a one-to-one mapping between the near and far field modes, which determines the radiation patterns, see Supplementary Material (SM) for details.
-Even though the radiation of subwavelength particles in free space via high-order modes is negligible, when there is one or more particles or excitation source at a distance from the particle, these modes play a more significant role in the radiation.
-In addition, we identify axial-eigenpermittivity sum
-rules, resulting in resonance splitting in anisotropic particles.
-We also show that the physics of anisotropic spheres is different from anisotropic ellipsoids: anisotropic spheres support uniaxial modes and exhibit eigenfrequency degeneracy, which leads to coherently superimposed modes, that anisotropic ellipsoids do not support.
-We apply our theory to spherical uniaxial and biaxial particles composed of hBN and MoO3 phonon supporting polar crystals, which can be synthesized, and present their unique spectra and eigenmodes [3, 64, 65, 66].
-Our starting point in calculating the eigenstates in the quasistatic regime is
-solving Laplace’s equation without a source, that reads for
-anisotropic media [67]
-where is an electric potential eigenstate that exists without a source for the eigenpermittivity tensor
-inside the inclusion and
-in the host medium ( is the identity matrix).
+While we consider the localized modes, in practice subwavelength particles radiate and there is a one-to-one mapping between the near and far field modes, which determines the radiation patterns, as we derive in Methods.
+Even though the radiation of subwavelength particles excited by a far field in free space via high-order modes is negligible, when there is one or more particles or excitation source at a distance from the particle, these modes play a significant role in the radiation.
+Results
+Our starting point in calculating the eigenstates of uniaxial and biaxial anisotropic spheres in the quasistatic regime is solving Laplace’s equation without a source, which reads for anisotropic media [67]
+where is an electric potential eigenstate that exists without a source for the eigenpermittivity tensor inside the inclusion and in the host medium ( is the identity matrix).
 Note that the physical permittivity tensor generally depends on and the resonances occur when it approximately satisfies the eigenpermittivity relations.
-Clearly, have to satisfy
-the boundary conditions at the particle interface.
+Clearly, have to satisfy the boundary conditions at the particle interface.
 Once the source-free Laplace’s equation and the boundary conditions are satisfied, these functions will be eigenstates or modes of the system by definition.
-In addition to the challenge that the symmetry of
-the material/crystal differs from the symmetry of the inclusion, Laplace’s equation has to be satisfied for anisotropic media, namely
-As a result, the widely-employed spherical harmonics are in general not eigenstates
-of anisotropic spherical particles.
-In Fig. 1, left column, we present isotropic, uniaxial, and biaxial particles, which are the most common types of particles.
-We now analyze the eigenmodes and eigenpermittivities of anisotropic
-spheres and ellipsoids. We start with the dipole response
-and proceed to the high-order mode responses.
-The anisotropic
-dipole eigenstates and eigenpermittivities of a sphere and ellipsoid
-can be deduced from the scattering analyses in the literature [42]
-and read for a dipole mode oriented along
-where in the case of an ellipsoid, are its intersections
-with the positive axes, respectively, and the second expression for the ellipsoid is on the external envelope. While these eigenstates
-and eigenpermittivities appear to be identical to the isotropic ones [36],
-here the modes are oriented along the resonant crystal axis independently
-of the incoming field polarization, even though it
-does affect the excitation magnitude.
-We now proceed to the high-order modes of anisotropic particles. In our analysis we choose to express the
-eigenstates in cartesian coordinates due to the crystal symmetry with functions
-that are suitable for anisotropic medium i.e., ones that enable when substituted in Laplace’s equation and satisfy boundary conditions.
-Since the field outside the sphere can be expanded with spherical harmonics,
-the high-order spherical harmonics are valid for isotropic
-medium [2, 35],
-and there
-is continuity of the potential on the particle envelope, we suggest
-that the anisotropic sphere modes are composed of specific isotropic-sphere modes or their superpositions [1, 2].
-To obtain eigenstates of a uniaxial sphere, we aim to arrive from anisotropic Laplace’s equation at without having to satisfy so that the modes can be excited when While the isotropic sphere eigenmodes and their superpositions satisfy anisotropic Laplace’s equation for when and it is sufficient to require to satisfy the equation.
+In addition to the challenge that the symmetry of the material/crystal differs from the symmetry of the inclusion, Laplace’s equation has to be satisfied for anisotropic media, namely As a result, the widely-employed spherical harmonics are in general not eigenstates of anisotropic spherical particles.
+In Fig.
+1, left column, we present isotropic, uniaxial, and biaxial spheres, which are the most common types of spheres.
+We now analyze the eigenmodes and eigenpermittivities of such anisotropic spheres.
+We start with the dipole response and proceed to the high-order mode responses.
+The anisotropic dipole eigenstates and eigenpermittivities of a sphere can be deduced from the scattering analyses in the literature [37] and read for a dipole mode oriented along
+While these eigenstates and eigenpermittivities appear to be identical to the isotropic ones [37], here the modes are oriented along the resonant crystal axis independently of the incoming field polarization, even though the polarization direction does affect the excitation magnitude.
+We now proceed to the high-order modes of anisotropic spheres.
+In our analysis we choose to express the eigenstates in cartesian coordinates due to the crystal symmetry with functions that are suitable for anisotropic medium i.e., ones that enable when substituted in Laplace’s equation and satisfy boundary conditions.
+Since the field outside the sphere can be expanded with spherical harmonics, the high-order spherical harmonics are valid for isotropic medium [31, 32], and there is continuity of the potential on the particle envelope, we suggest that the anisotropic sphere modes are composed of specific isotropic-sphere modes or their superpositions [41, 31].
+To obtain eigenstates of a uniaxial sphere, we aim to arrive from anisotropic Laplace’s equation at without having to satisfy so that the modes can be excited when While the isotropic sphere eigenmodes and their superpositions satisfy anisotropic Laplace’s equation for when and they naturally satisfy the above requirement.
 Interestingly, such functions can be composed of the isotropic-sphere eigenstates Thus, we find the uniaxial sphere modes and obtain their eigenpermittivity relations from the boundary conditions to arbitrary order:
-It can be seen that since in the boundary conditions and modes composed of and satisfy the isotropic eigenpermittivities [1, 2] for the equal permittivity axes and new eigenpermittivity relations, respectively.
+where correspond to and the first condition holds if either or , whereas the second holds only for resulting in resonance splitting for (two closely-spaced resonances instead of one for isotropic spheres).
+One can see that since in the boundary conditions and modes composed of and satisfy the isotropic eigenpermittivities [41, 31] for the equal permittivity axes and new eigenpermittivity relations, respectively.
 However, our modes in both cases are directional as opposed to the isotropic sphere modes.
-Alternatively, one could select the basis .
+Alternatively, one could select the basis for uniaxial spheres.
 Note also that the resonance conditions depend on which can be utilized to sense the environment in which the particle is situated.
-Figure 2: Spectrum and eigenmodes of a subwavelength uniaxial sphere.
-(a) A uniaxial sphere and the permittivity of bulk hBN.
-(b) The scattering spectrum and of the dipole modes on an hBN sphere surface.
-Here we see scattering peaks at approximately (red line) and (blue line) associated with dipole modes in and respectively.
-The crystal axes [100],[010],[001] are directed along the axes, respectively.
-(c) The scattering spectrum and of the second-order modes, where the and modes and the uniaxial and modes are both doubly degenerate in frequency and there is a resonance splitting compared to the bulk material.
-(d) Scattering spectrum and of the third-order mode, where is excited at two s.
-Moreover, there are geometric-anisotropic shifts compared to the bulk for all modes.
-Interestingly, high-order resonances have lower compared to the dipole modes, unlike isotropic spheres.
-In some cases superpositions of the isotropic sphere modes satisfy and anisotropic Laplace’s equation is valid for arbitrary values of , which have to obey only the boundary condition.
-Thus, we suggest the following biaxial-sphere modes and calculate the eigenpermittivity relations from the boundary condition:
-Note that this class of functions preserves the form under Thus, the axial permittivities satisfy novel and interesting sum rules,
-this time for a general biaxial medium.
+In some cases, superpositions of the isotropic sphere modes satisfy and anisotropic Laplace’s equation is valid for arbitrary values of , which have to obey only the boundary condition.
+Thus, we derive the following biaxial-sphere modes and calculate the eigenpermittivity relations from the boundary condition:
+These eigenstates are superpositions of spherical harmonics in all space, as is explained in Methods.
+Note that when imposing the boundary condition, this class of functions preserves the functional form under inside the particle.
+Thus, the axial permittivities satisfy novel and interesting sum rules, this time for a general biaxial medium.
+These sum rules result in resonance splitting compared to the isotropic-sphere resonances: anisotropic bulk media usually exhibit a distinct resonance in each axis.
+Thus, for a resonance in a given axis in the bulk medium (), usually the sums result in two resonances occurring at different frequencies.
+Hence, for a biaxial material whose bulk exhibits three resonances, our resonance conditions predict 18 different peaks in the spectrum in addition to the 3 dipole resonances.
+In the Supplementary Material (SM) we show that the derived uniaxial modes form a complete set and if additional biaxial modes exist they are typically negligible.
 In Fig.
-1 we present the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles.
-It can be seen that anisotropic particles support significantly different modes, which appear in several spectral regions, and their direction is determined only by the crystal orientation.
+1 we present a scheme of the angular distributions of the electroquasistatic modes of isotropic and anisotropic particles.
+One can see that anisotropic particles support significantly different modes, which typically appear in several spectral regions as each axial permittivity experiences a resonanance at a different frequency (colors correspond to frequencies).
+In addition, their direction is determined only by the crystal orientation.
 Moreover, anisotropic particles exhibit rich physics with resonance splitting for both uniaxial and biaxial particles and degeneracy in frequency for uniaxial particles, which leads to superimposed modes.
-Figure 3: Spectrum and eigenmodes of a subwavelength biaxial particle. (a) A biaxial sphere
-with the bulk permittivities of .
-The scattering spectra and of the eigenmodes on a sphere surface for the dipole (b), second-order (c), and third-order (d) modes.
-Note that there are geometric-anisotropic frequency shifts for all modes and resonance splitting for the second-order modes compared to the bulk.
-Furthermore, for a biaxial material whose bulk exhibits three resonances, our resonance conditions predict 18 different peaks in the spectrum in addition to the 3 dipole resonances.
-In the SM we show that the derived uniaxial modes are all the modes and if additional biaxial modes exist they are typically negligible.
 In Fig.
@@
 In Fig.
+2 (e) we show in the radiation patterns of the uniaxial sphere excited by an oscillating dipole (Methods).
+In Fig.
 3 we present a biaxial sphere and the axial permittivities of bulk with three resonances.
@@
 Similarly to the uniaxial particle, the second-order modes exhibit resonance splitting and all the modes display geometric-anisotropic frequency shifts.
-However, here the degeneracy is lifted since there are three different permittivities and the uniaxial modes are not present.
-Interestingly, in both uniaxial and biaxial spheres, the high-order mode resonances can have lower frequencies compared to the dipole mode, unlike the situation in isotropic spheres.
-We now consider the high-order modes of anisotropic ellipsoids in analogy to the anisotropic sphere.
-Interestingly, the biaxial-sphere modes inside the inclusion volume correspond to the second type of Niven’s functions used in isotropic ellipsoids, at least up to fourth order [36, 68, 69, 70]
-We utilize inside the anisotropic
-ellipsoid, similarly to the isotropic case of an ellipsoid in Ref. [36].
-To impose continuity of for we project in the
-direction perpendicular to the interface with the unit vector
-Comparing inside the ellipsoid to the isotropic case
-we obtain that since and (inside and outside) are the same, we can equate to get an eigenmode and obtain the general sum rule
-and for the sphere limit we get as expected.
-The modes in the form similarly to the uniaxial sphere modes, have eigenpermittivity relations that are too restrictive and not likely to be realized experimentally.
-We proceed to the third-order mode of an anisotropic ellipsoid. We write the potential and field inside the inclusion
-Similarly, we compare inside the ellipsoid in the isotropic and anisotropic cases and obtain that a resonance occurs when
-We take the anisotropic-sphere limit of the eigenpermittivity relation and obtain agreement
-In this way one can also find the fourth-order biaxial ellipsoid mode and resonance condition.
-Finally, we verified our theory for uniaxial and biaxial spheres and a biaxial ellipsoid in simulations (see Appendix) and also experimentally for two biaxial ellipsoids with hyperspectral midIR near-field imaging in Ref.
-[56], both with excellent agreement for the modes and the spectra.
-In conclusion, we presented a framework to derive the
-modes and eigenpermittivity relations of anisotropic spherical and
-ellipsoid particles in the quasistatic regime. While isotropic structures
-have discrete eigenpermittivities, we showed that anisotropic particles exhibit eigenpermittivity sum rules that couple
-the axial permittivities.
-We identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes, and resonance splitting in anisotropic particles.
-Our analysis revealed that both particle types support biaxial modes but anisotropic nanospheres support uniaxial modes that anisotropic ellipsoids typically do not support.
-Importantly, our derived modes can be used to calculate the eigenstates and resonance conditions of clusters of anisotropic particles [35].
-The multispectral behavior of anisotropic particles may allow them to function as a biomarker with enhanced specificity of the spectral signature.
-In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection [60].
-Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [61, 71, 72].
-Furthermore, our derived modes of anisotropic particles are expected to have implications for the radiation patterns in quantum transitions of atoms in magnetic fields and anisotropic molecules.
-Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing.
-Finally, our results
-apply also to other fields of physics such as heat conduction and quasimagnetostatics [57].
-Acknowledgement
-E. E. Narimanov, A. Niv, Y. Mazor, J. Jeffet, and Y. Reches are acknowledged for the insightful comments.
-We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.
-I Appendix
-Comparison of the theory to simulations
-Figure 4: Scattering spectrum, peak widths, and fields near resonance for a uniaxial hBN sphere of a radius excited by an oscillating dipole located at with a dipole moment of .
-Top: over the particle envelope calculated in COMSOL and compared to the peaks predicted analytically using with excellent agreement.
-The circles are the peak widths obtained analytically from compared to the ones calculated from the simulations, with very good agreement.
-(a),(b),(c),(d),(e) for the resonance conditions: calculated at
-respectively, using COMSOL compared to the mode fields calculated analytically with very good agreement.
-Bottom: over the particle envelope calculated in COMSOL for a dipole located at (same dipole distance as the previous case) with a dipole moment of As predicted the modes disappeared from the spectrum.Figure 5: Scattering spectrum and modes/fields near resonance for a biaxial -MoO3 sphere of a radius excited by an oscillating dipole located at with a dipole moment of .
-Top: calculated in COMSOL and compared to the peaks predicted analytically using with excellent agreement.
-(a),(b),(c),(d),(e),(f),(g) for the resonance conditions: calculated at respectively, using COMSOL compared to the mode fields calculated analytically with very good agreement.
-Figure 6: Scattering spectrum and modes/fields near resonance for a biaxial -MoO3 ellipsoid with excited by an oscillating dipole located at with a dipole moment of .
-Top: Surface integral over the particle envelope calculated in COMSOL and compared to the peaks predicted analytically using with excellent agreement.
-(a),(b),(c),(d),(e),(f),(g) for the resonance conditions: calculated at respectively, using COMSOL with very good agreement to the mode fields calculated analytically (same as in Fig.
-5, defined in the ellipsoid volume).
+However, here the degeneracy is lifted since there are three different permittivities and the purely-uniaxial modes are not present.
+Interestingly, in both uniaxial and biaxial spheres, the high-order mode resonances can have lower frequencies compared to the dipole mode, in stark contrast to the situation in isotropic spheres.
+Fig.
+3 (e) presents in the radiation patterns of representative biaxial-sphere modes.
+We now analyze the high-order modes of uniaxial and biaxial ellipsoids in analogy to the anisotropic spheres.
+Similarly, we suggest that the anisotropic ellipsoid modes are specific modes or superpositions of isotropic ellipsoid modes and consider the modes expressed in cartesian coordinates (Niven’s functions).
+We apply our conditions with the second derivatives to these modes and observe that while the first type of Niven’s functions do not satisfy them, the second type do satisfy them and inside the inclusion correspond to the biaxial sphere modes [68, 69, 70, 71].
+To impose continuity of for we project in the direction perpendicular to the interface with the unit vector and obtain for
+Comparing inside the ellipsoid to the isotropic case we obtain that since and (inside and outside) are the same for this isotropic and anisotropic cases, we can equate to obtain the general (sum rule) resonance condition:
+where [68]
+For the sphere limit we get as expected.
+The modes in the form similarly to the uniaxial sphere modes, have eigenpermittivity relations that are too restrictive (two conditions instead of one) and not likely to be realized experimentally.
+We proceed to the third-order mode of an anisotropic ellipsoid.
+We write the potential and field inside the inclusion Similarly, we equate inside the ellipsoid in the isotropic and anisotropic cases and obtain the resonance condition:
+We take the anisotropic-sphere limit of the eigenpermittivity relation and obtain agreement In this way one can readily find the fourth-order biaxial ellipsoid mode and resonance condition.
+Finally, we demonstrate our theory for spheres and ellipsoids composed of hBN and in full-wave electrodynamic simulations.
+While our theory applies to any uniaxial and biaxial material, we have chosen these materials as they are currently at the forefront of photonics research and enable to achieve high Q factors.
 We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes and the permittivity model used in Figs.
 2,3 [54].
-The sphere is of radius and we situate an oscillating dipole at with a dipole moment of We perform frequency domain simulations with a frequency sweep from (1/s) to (1/s) with 304 frequencies using extensive COMSOL simulations.
-The dipole moment was chosen to be in the direction of since we assume [67, 73, 2], and later on prove, that the mode excitation coefficient depends on so that we need to incorporate only preserving the angular dependence in .
-This, together with the fact that the dipole has only components, ensures that the degenerate modes will add constructively, see SM for details.
-Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflects the presence of higher spatial frequencies near the dipole, which excite higher-order modes.
-We then calculate the peak widths analytically and from the simulations.
-To demonstrate the mode coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the axis.
-In this case we do not expect the modes to show in the spectrum since in order to excite them a component of the dipole location is required.
+The sphere has a radius of and we situate an oscillating dipole at with a dipole moment of We perform frequency domain simulations with a frequency sweep from (1/s) to (1/s) with 304 frequencies using extensive COMSOL simulations.
+The dipole moment was chosen to be in the direction of since we assume [67, 72, 31], and prove in the SM, that the mode excitation coefficient depends on so that we need to incorporate only preserving the angular dependence in .
+This, together with the fact that the dipole has only components, ensures that the degenerate modes will add constructively, as is explained in the Methods Section.
+Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflect the presence of higher spatial frequencies near the dipole, which excite higher-order modes.
+For a general incoming field, the mode excitation coefficient is given by where is a window function that equals one in the particle volume, and we utilized a dipole source since a far field excites high-order modes with negligible magnitudes; more details are given in the SM.
+We then calculate the resonance peak widths analytically (from 1/resonance condition) and from the simulations.
+To demonstrate the mode-coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the axis.
+In this case, we do not expect the modes to show in the spectrum since in order to excite them a component of the dipole location is required.
 In Fig.
-4 we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of compared to the surface integral of over the uniaxial particle envelope calculated in COMSOL with excellent agreement.
-Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows in Fig.
-4) and mode degeneracy for uniaxial spheres.
-We also obtained the peak widths analytically from and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to is altered due to the peak on the right, and therefore we haven’t accounted for it).
-Clearly, there is remarkable agreement between them.
-This showcases our important ability to analytically predict the Q factors given by
-We then plot for all the observed peaks inside the particle calculated in COMSOL and obtained analytically from the modes.
+4 (top panel) we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of compared to the surface integral of over the uniaxial particle envelope calculated in COMSOL with excellent agreement.
+Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows) and mode degeneracy for uniaxial spheres.
+We also obtained the peak widths analytically from and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to is altered due to the peak on the right, and therefore we haven’t analyzed it).
+Clearly, there is remarkable agreement between them, which showcases our important ability to analytically predict the Q factors given by We then plot for all the observed peaks inside the particle calculated in COMSOL and obtained analytically from the modes (Methods).
 While the COMSOL fields include the incoming field from the oscillating dipole, close to a resonance this is typically negligible and indeed we observe very good agreement.
@@
 This highlights the directional nature of the modes as opposed to isotropic particles.
-We then considered a biaxial subwavelength sphere composed of MoO3 with the permittivity model in Ref. [55]. The sphere is excited by a dipole located at with the dipole moment of
-We performed frequency domain simulations with a frequency sweep between and with 175 frequencies using COMSOL.
+We then considered a biaxial subwavelength sphere composed of MoO3 with the permittivity model in Ref.
+[54].
+The sphere is excited by a dipole located at with the dipole moment of We performed frequency domain simulations with a frequency sweep between and with 175 frequencies using COMSOL.
 Here, again, the dipole moment is directed along to preserve the angular dependency of the mode in the mode-excitation coefficient.
@@
 [54].
-The ellipsoid is excited by a dipole located at with the dipole moment of
-We performed frequency domain simulations with a frequency sweep between and with 175 frequencies using COMSOL.
+The ellipsoid is excited by a dipole located at with the dipole moment of We performed frequency domain simulations with a frequency sweep between and with 175 frequencies using COMSOL.
 In Fig.
 6 we present the result of the surface integral over the particle envelope from the COMSOL simulations and the predicted peaks from the resonance conditions with very good agreement.
-The predicted peaks were easily calculated by finding the maxima of .
-We then compared for all the observed resonances from the simulations to the ones obtained analytically from the modes with excellent agreement.
-We note that the isotropic ellipsoid eigenpermittivities that were calculated to obtain the biaxial ellipsoid resonance permittivity relations are the following:
-As can be seen, for an isotropic ellipsoid with these semi axis ratios and and due to the anisotropic material these resonance are split, as can be seen in Fig.
+The predicted peaks were readily calculated by finding the maxima of .
+We then compared for all the observed resonances from the simulations to the ones obtained analytically from the modes (equivalent to Fig.
+5) with excellent agreement.
+We note that the isotropic-ellipsoid eigenpermittivities that were calculated to obtain the biaxial ellipsoid resonance permittivity relations are the following: As can be seen, for an isotropic ellipsoid with these semi axis ratios and and due to the anisotropic material these resonances are split, as can be seen in Fig.
 6.
-This differs from the case of spheres where the splitting occurs for modes of the same order.
-Also, in this case the anisotropy of the material could in principle lead to resonance merging.
-Finally, there is a clear physical intuition for the resonance conditions of anisotropic ellipsoids: the resonances of isotropic ellispoids, on which they are based, tend to favor shorter ellipsoid axes, since they exhibit higher real permittivity values that have smaller imaginary parts of the physical permittivity, leading to stronger and sharper resonances.
-II Supplementary Material
-II.1 Correspondence between the derived modes and the radiation patterns
+In contrast to spheres, where the degeneracy is enforced by symmetry, here the degeneracy is accidental.
+Finally, there is a clear physical intuition for the resonance conditions of anisotropic ellipsoids: the resonances of isotropic ellipsoids, on which they are based, tend to favor shorter ellipsoid axes, since they exhibit higher real permittivity values that have smaller imaginary parts of the physical permittivity, leading to stronger and sharper resonances.
+Discussion
+We derived the modes and eigenpermittivity relations of uniaxial and biaxial particles in the quasistatic regime.
+We showed that while isotropic spheres have discrete eigenpermittivities, anisotropic spheres exhibit eigenpermittivity sum rules that couple the axial permittivities, leading to resonance splitting.
+We also identified eigenfrequency degeneracy for uniaxial spheres, which leads to superimposed modes.
+We generalized our results to uniaxial and biaxial ellipsoids, with eigenpermittivity sum rules that account for both the anisotropic material properties and the ellipsoid’s geometry.
+Our derived modes and resonance conditions have implications for the thermal emission, enhancement of spontaneous emission, and energy transfer of a single anisotropic particle, as well as for the response of clusters of anisotropic particles [73, 74, 75, 76, 32].
+The multispectral behavior of uniaxial and biaxial particles may allow them to function as biomarkers with enhanced specificity of the spectral signature.
+In addition, the environment-dependent resonance shifts of the particle can be harnessed to measure the biological environment, which may lead to improved cancer detection [59].
+Moreover, due to the directional nature of the modes in anisotropic particles, another potential application of our resonators is detection of dark matter [60, 77, 78].
+Furthermore, as our derived modes of anisotropic particles correspond to novel radiation patterns, they are expected to have implications for the emission in quantum transitions of atoms in magnetic fields and anisotropic molecules.
+Future research could focus on analyzing their nonlinear optical response with potential applications in optical computing.
+Finally, our results apply also to other fields of physics such as heat conduction and quasimagnetostatics [56].
+Acknowledgement
+E.
+E.
+Narimanov, A.
+Niv, Y.
+Mazor, J.
+Jeffet, and Y.
+Reches are acknowledged for the insightful comments.
+We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.
+I Methods
+I.1 Correspondence between the derived modes and the radiation patterns
 Here we derive the correspondence between the localized modes that we obtained in the main text and the radiation patterns of the anisotropic subwavelength spheres.
-We utilize the TM vector spherical harmonics (VSH).
-This field pattern is proportional to the sphere mode outside the inclusion [74], which gives the quasistatic results in the limit of Even though this mode is usually considered for isotropic spheres, the results can be applied to the modes of anisotropic spheres since their modes are superpositions of spherical harmonics, which correspond to superpositions of VSHs.
-We write a TM VSH:
-where is the spherical Hankel function of the first kind.
-Using the vector identity for
-we get for the first term:
-where
-For the second term we use the identity
-to get
-Then, we utilize the spherical-harmonic identity
+We utilize the TM vector spherical harmonics (VSH) modes for a general spherical particle given by where
+and are the spherical bessel function and spherical Hankel function of the first kind and [79, 20].
+Calculating the eigenpermittivity from these modes in the limit of gives the quasistatic eigenvalues.
+For example, the full electrodynamic mode gives the following eigenpermittivities for a sphere with nm and nm : which are very close to the quasistatic results of Even though these modes are usually considered for isotropic spheres, the results can be applied to anisotropic spheres since their modes are superpositions of spherical harmonics, which correspond to superpositions of VSHs.
+We write a TM VSH outside the sphere:
+Using the vector identity for we get for the first term:
+For the second term we employ the same identity to get
+Then, we utilize the spherical-harmonic identity [80]
 to obtain
 All in all we get:
-We see that when the second term, which is longitudinal, dominates with scaling, and for
-the first term, which is transverse, dominates (actually its first term that scales as
-Thus, one can conclude from the superposition of that participate in the
-quasistatic anisotropic modes, that the radiation pattern will be of the form:
-Please note that which gives
-overall scaling of the radiated field and we get the following radiation pattern:
-II.2 Biaxial sphere modes as superpositions of spherical harmonics
+We see that when since the overall scaling is of in agreement with the quasistatic analysis.
+For since and the first term, which is transverse and has scaling as expected, dominates.
+Thus, one can conclude from the superposition of that participate in the quasistatic anisotropic modes, that the radiation pattern will be of the form:
+where the sum is over the mode orders which participate in the anisotropic-sphere modes. Please note that which gives overall scaling of the radiated field and we get the following radiation pattern:
+One can readily see that the mode of a subwavelength sphere gives the dipole radiation pattern of .
+In addition, the radiation patterns of high-order anisotropic particle modes differ from the isotropic case.
+Note that while the uniaxial-sphere modes are identical to the isotropic sphere modes, in nearly all practical scenarios, their radiation patterns differ due to the fact that an excitation source will excite distinct superpositions of modes in the two cases.
+I.2 Biaxial sphere modes as superpositions of spherical harmonics
 We express the second-order biaxial modes as superpositions of spherical harmonics:
@@
 Note that the modes and their resonance conditions readily follow.
-II.3 Completeness of the solution
+I.3 Superimposing the degenerate uniaxial modes for the sphere simulation comparison
+I.3.1 Theory
+We now analytically derive the field distributions and resonance frequencies for the uniaxial sphere and the dipole location in the main text.
+Dipole mode
+The resonance condition is and we therefore find the maximum of which occurs for hBN [54] at The mode field inside the particle is given by
+2nd order modes
+For the first 2nd-order mode we get from the theory which for hBN approximately occurs at We add the degenerate modes, which are excited with equal strengths:
+second mode
+The derived resonance condition reads
+For hBN this is approximately satisfied for Now we calculate the field by adding the degenerate modes:
+3rd order modes
+First mode
+The resonance condition reads:
+For hBN this occurs at . Adding the degenerate modes we get
+To simplify things, we look at the plane
+Second mode
+The resonance condition is satisfied for hBN for Adding the modes we obtain:
+We plot these field magnitude distributions in Fig. 4.
+One can express the eigenvalue in the denominator in terms of the physical parameter using the eigenpermittivity relations, obtaining exactly the resonance conditions now for the physical parameters used along the paper (e.g., ones used in Figs.
+2 and 3).
+Note that the first sum includes only the dipole term and the second sum includes the remaining eigenstates in agreement with our analysis along the paper.
+It is worth noting that this formalism should also apply to other types of anisotropy such as spherical.
+Supplementary Material
+Completeness of the solution
 We now prove that the derived uniaxial-sphere modes are all the eigenmodes.
@@
 Therefore, we expect that the dominant biaxial modes satisfy and considered all such possible modes in the main text.
-II.4 Superimposing the degenerate uniaxial modes for the sphere simulation comparison
-II.4.1 Theory
-We now analytically derive the field distributions and resonance frequencies for the uniaxial sphere and the dipole location in the Appendix.
-Dipole mode
-The resonance condition is and we therefore find the maximum of
-which occurs for hBN [55] at The mode field inside the particle is given by
-2nd order modes
-For the first 2nd-order mode we get from the theory
-which for hBN approximately occurs at We add the degenerate modes, which are excited with equal strengths:
-second mode
-The derived resonance condition reads
-For hBN this is approximately satisfied for
-Now we calculate the field by adding the degenerate modes:
-3rd order modes
-First mode
-The resonance condition reads:
-For hBN this occurs at
-. Adding the degenerate modes we get
-To simplify things, we look at the plane
-Second mode
-The resonance condition is satisfied
-for hBN for Adding the modes we obtain:
-We plot these field magnitude distributions in Fig. 4.
-II.5 Eigenmode expansion for an anisotropic inclusion
+Eigenmode expansion for an anisotropic inclusion
 Eigenstate expansions have been derived for an isotropic inclusion and for an inclusion with only one axial permittivity that differs from the host medium permittivity [79, 67].
 Here we expand the quasielectrostatic potential for an anisotropic inclusion.
-For concreteness, we consider a uniaxial inclusion and the derived eigenstates. Generalizing it to a biaxial inclusion readily follows.
+For concreteness, we consider a uniaxial inclusion and the derived eigenstates.
+Generalizing it to a biaxial inclusion readily follows.
 We start by writing Poisson’s equation for a medium with an anisotropic inclusion:
 where is a step function that equals 1 inside the inclusion, is the inclusion permittivity tensor, and is the host medium permittivity.
-Now we move one term to the lhs in order to be able to define an
-eigenvalue equation.
-where . The Green function in this case will be for a medium with an inclusion
-with anisotropy in the axis and axial permittivities that are equal to the host medium permittivity in the other axes.
+Now we move one term to the lhs in order to be able to define an eigenvalue equation.
+where . The Green function in this case will be for a medium with an inclusion with anisotropy in the axis and axial permittivities that are equal to the host medium permittivity in the other axes.
 where we defined to adjust it to an eigenvalue equation with an operator operating on and we performed integration by parts.
-We express the electric potential and expand it using the anisotropic
-eigenfunctions defined by Laplace’s equation. In the following is defined for the particle with anisotropy in the axis. We write:
+We express the electric potential and expand it using the anisotropic eigenfunctions defined by Laplace’s equation. In the following is defined for the particle with anisotropy in the axis. We write:
 Now we expand using the eigenstates of a particle with anisotropy in the axis:
-where is the electric potential for the dipole in free space.
-All in all we get:
-One can express the eigenvalue in the denominator in terms of the physical parameter using the eigenpermittivity relations, obtaining exactly the resonance conditions now for the physical parameters used along the paper (e.g., ones used in Figs.
-2 and 3).
-Note that the first sum includes only the dipole term and the second sum includes the remaining eigenstates in agreement with our analysis along the paper.
-It is worth noting that this formalism should also apply to other types of anisotropy such as spherical.
-Note that in Fig.
-4 there is a very small peak at This peak may be due to an off-resonance uniaxial mode as could be understood from the above expansion when applied to biaxial inclusions.
-Such an expansion is composed of three sums where the second one corresponds to the off resonance uniaxial-mode contributions.
-This qualitatively agrees with the potential calculated from the field in the simulation results of the form similarly to the mode constructed by adding the modes.
-Interestingly, this very small peak is not present in the spectrum of the biaxial ellipsoid, in agreement with our prediction that ellipsoids do not favor uniaxial modes.
+where is the electric potential for the dipole in free space. All in all we get:

### v6

@@
 Anisotropic subwavelength particles—ubiquitous in natural and engineered systems—offer new degrees of freedom that couple geometry and material properties, unlocking previously inaccessible spectral regions.
-Here, we establish a universal quasistatic framework describing the eigenmodes and resonance conditions of uniaxial and biaxial nanoparticles.
-Closed-form solutions reveal axial-permittivity sum rules and a fundamental material-anisotropy-induced symmetry breaking, manifesting as resonance splitting and novel radiation patterns.
+Here, we establish a universal full-wave framework describing the eigenmodes and resonance conditions of uniaxial and biaxial nanoparticles.
+Closed-form solutions reveal axial permittivity sum rules and a fundamental material-anisotropy-induced symmetry breaking, manifesting as resonance splitting and novel radiation patterns.
 Generalizing the theory to ellipsoids provides geometric tunability of the multispectral response, while analytic predictions of quality factors elucidate how anisotropy governs mode localization and energy loss.
-Full-wave simulations of h-BN and -MoO3 nanoparticles confirm the theory and highlight its predictive power.
+Full-wave simulations of h-BN and -MoO3 nanoparticles, together with our recently reported experimental observations, confirm the theory and highlight its predictive power.
 This framework unifies the understanding of anisotropic nanostructures across optics, thermal transport, and magnetism, opening pathways to a new generation of photonic devices with tunable multispectral response and controlled emission patterns, including directional detectors and spectrally selective mid infrared sensors.
@@
 We take the anisotropic-sphere limit of the eigenpermittivity relation and obtain agreement In this way one can readily find the fourth-order biaxial ellipsoid mode and resonance condition.
+| (a) | (b) |
 Finally, we demonstrate our theory for spheres and ellipsoids composed of hBN and in full-wave electrodynamic simulations.
@@
 This highlights the directional nature of the modes as opposed to isotropic particles.
+In Fig.
+5 we compared the radiation pattern from the full-wave closed-form solution corresponding to the mode with the resonance condition outside a uniaxial particle to a full-wave COMSOL simulation of the scattered field of a uniaxial particle with nm and excited by a current source similarly to Fig.
+4 (top) in the radiation zone in the radiation zone .
+Note that the COMSOL simulations utilized the axial symmetry and are therefore presented on half a sphere.
+As can be seen, the radiation-pattern results show very good agreement.
 We then considered a biaxial subwavelength sphere composed of MoO3 with the permittivity model in Ref.
@@
 Crucially, our simulations confirm the resonance splitting effect predicted analytically, see arrows in Fig.
-5.
+6.
 We then compared for all the observed resonances from the simulations to the ones obtained analytically from the modes with excellent agreement.
@@
 We then compared for all the observed resonances from the simulations to the ones obtained analytically from the modes (equivalent to Fig.
-5) with excellent agreement.
+6) with excellent agreement.
 We note that the isotropic-ellipsoid eigenpermittivities that were calculated to obtain the biaxial ellipsoid resonance permittivity relations are the following: As can be seen, for an isotropic ellipsoid with these semi axis ratios and and due to the anisotropic material these resonances are split, as can be seen in Fig.
-6.
+7.
 In contrast to spheres, where the degeneracy is enforced by symmetry, here the degeneracy is accidental.
@@
 Mazor, J.
-Jeffet, and Y.
-Reches are acknowledged for the insightful comments.
+Jeffet, Y.
+Reches, and I.
+Kalyan are acknowledged for the insightful comments.
 We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.
@@
 Using the vector identity for we get for the first term:
-For the second term we employ the same identity to get
-Then, we utilize the spherical-harmonic identity [80]
+For the second term we arrive at
+where we have used the vector identity for Then, we utilize the spherical-harmonic identity [80]
 to obtain
-All in all we get:
+All in all, we get:
 We see that when since the overall scaling is of in agreement with the quasistatic analysis.
 For since and the first term, which is transverse and has scaling as expected, dominates.
-Thus, one can conclude from the superposition of that participate in the quasistatic anisotropic modes, that the radiation pattern will be of the form:
+Thus, one can conclude from the superposition of that participate in the quasistatic anisotropic modes, that the radiation patterns are of the form:
 where the sum is over the mode orders which participate in the anisotropic-sphere modes. Please note that which gives overall scaling of the radiated field and we get the following radiation pattern:
@@
 It is worth noting that this formalism should also apply to other types of anisotropy such as spherical.
-Supplementary Material
-Completeness of the solution
-We now prove that the derived uniaxial-sphere modes are all the eigenmodes.
-Let us assume that there is an additional eigenstate besides the four modes for each .
-Since the isotropic-sphere modes span the space inside the sphere, this eigenstate is a superposition of the remaining isotropic-sphere modes
-An eigenstate must satisfy anisotropic Laplace’s equation inside the inclusion, which reads:
-Since it is a superposition of isotropic-sphere modes we write:
-However, all these modes have orders of that are 2 or higher (including from ) and each function is fundamentally different from the others.
-Hence, such an eigenstate does not exist, and we have a contradiction.
-Conclusion: the derived uniaxial sphere modes are all the modes.
-Note that the uniaxial mode basis does not span space as in the standard case.
-Similarly, we write for a biaxial particle
-If we get which is a uniaxial medium, and we therefore exclude this case.
-All the modes are isotropic (from the boundary conditions) and the modes satisfy when Even if such modes exist, their axial permittivities must satisfy both Laplace’s equation (in a non-trivial manner) and the boundary condition, which imposes strong constraints.
-As a result, their potential contribution is typically weak.
-Therefore, we expect that the dominant biaxial modes satisfy and considered all such possible modes in the main text.
-Eigenmode expansion for an anisotropic inclusion
-Eigenstate expansions have been derived for an isotropic inclusion and for an inclusion with only one axial permittivity that differs from the host medium permittivity [79, 67].
-Here we expand the quasielectrostatic potential for an anisotropic inclusion.
-For concreteness, we consider a uniaxial inclusion and the derived eigenstates.
-Generalizing it to a biaxial inclusion readily follows.
-We start by writing Poisson’s equation for a medium with an anisotropic inclusion:
-where is a step function that equals 1 inside the inclusion, is the inclusion permittivity tensor, and is the host medium permittivity.
-Now we move one term to the lhs in order to be able to define an eigenvalue equation.
-where . The Green function in this case will be for a medium with an inclusion with anisotropy in the axis and axial permittivities that are equal to the host medium permittivity in the other axes.
-where we defined to adjust it to an eigenvalue equation with an operator operating on and we performed integration by parts.
-We express the electric potential and expand it using the anisotropic eigenfunctions defined by Laplace’s equation. In the following is defined for the particle with anisotropy in the axis. We write:
-Now we expand using the eigenstates of a particle with anisotropy in the axis:
-where is the electric potential for the dipole in free space. All in all we get: